# Scale dependence improvement of the quartic scalar field thermal effective potential in the optimized perturbation theory

### v1

@@ V Results
Let us now analyze how well the VRG performs in order to mitigate the scale dependence of the OPT effective potential by mainly comparing predictions for the symmetric ($m^{2}\geq 0$) and the non-symmetric ($m^{2}<0$) cases. Concerning the first case, we will also make a comparison with the results provided by other nonperturbative methods commonly used in the literature, e.g., the results obtained from the 2-particle irreducible approach (2PI), from the functional renormalization group (FRG) and from the RGOPT method. This will help us gauge how our approach performs compared to these other methods. In addition, we also examine whether the VRG potential preserves the universality class of the $\lambda\phi^{4}$ theory [^1] Recall that the $Z_{2}$ symmetric scalar field model (1) belongs to the same universality class of the Ising model for $d\leq 4$, with the model exhibiting a second-order phase transition at the critical point [70, 59, 71]..
@@
To analyze the dependence of the effective potential on the renormalization scale, we start by considering the scale dependence of the renormalized parameters $m(\mu)^{2}$, $\lambda(\mu)$, $\varphi(\mu)$, and $\Lambda(\mu)$. As already specified in the previous section, at the first order for both OPT and VRG, we consider the running of the renormalized parameters up to ${\cal O}(\hbar^{2})$ and which are determined by the equations:
@@
$\mu\frac{\Lambda^{\prime}(\mu)}{\Lambda(\mu)}$ $=$ $\frac{\lambda(\mu)}{2(4\pi)^{2}},$ (29)
@@
$\mu\frac{\Lambda^{\prime}(\mu)}{\Lambda(\mu)}$ $=$ $\frac{\lambda(\mu)}{2(4\pi)^{2}}+\frac{\lambda(\mu)^{3}}{16(4\pi)^{6}},$ (33)
where we now substitute these solutions in Eq. (6), up to order $\delta^{2}$, in the case of OPT, and in Eq. (72) for VRG. The boundary conditions used to solve the set of Eqs. (26)-(33) are: $\lambda(\mu_{0})=\lambda_{0}$, $m(\mu_{0})=m_{0}$, $\varphi(\mu_{0})=\varphi_{0}$, and $\Lambda(\mu_{0})=\Lambda_{0}$ where $\mu_{0}$ is a reference scale to be defined below. As is common in the literature, to check the stability of the computed quantities with the scale $\mu$, we will vary it around the so-called “central” value $\mu=2\pi T$ within the usual range $\mu\in[\pi T,4\pi T]$. As a reference scale, we can then choose $\mu_{0}=2\pi T_{0}$ where, for simplicity, the reference temperature is set to $T_{0}=m_{0}/(2\pi)$ so that $\mu_{0}=m_{0}$. This choice will allow us to easily express all physical quantities in units of $m_{0}$ (or likewise, in terms of the reference scale $\mu_{0}$).
V.1 Symmetric Phase
@@ V.2 Broken Phase
As usual, the order parameter $\sigma(T)$ can be determined by minimizing the thermal effective potential,
@@
Figure 8: Subtracted effective potential, $\Delta V=V_{\rm eff}(\varphi,T)-V_{\rm eff}(\varphi=0,T)$, as a function of $\varphi$. In both cases, the results for OPT and VRG are at order $\delta^{2}$. The coupling is fixed at the values $\lambda_{0}=0.1$ (panel a) and $\lambda_{0}=1$ (panel b). The scale dependence range is such that $\pi T\leq\mu\leq 4\pi T$. From bottom to top, the regions correspond to the temperature fixed at $T_{c}-\Delta T, T=T$$T_{c}$, and $T=T_{c}+\Delta T$, respectively, with $\Delta T/m_{0}=0.01$.
@@
Let us now check the dependence of the effective potential in the OPT and VRG cases as a function of the background field just as we have done in the symmetric case (see Fig. 5). This is shown in Fig. 8. Two values of the coupling constant have been chosen, $\lambda_{0}=0.1$ in Fig. 8(a) and $\lambda_{0}=1$ in Fig. 8(b). We have also chosen values for the temperature below, at and above the critical value, which illustrates well that the phase transition is second-order, as also already confirmed with the results shown in Fig. 6. Furthermore, the results displayed in Fig. 8 indicate the efficiency of the VRG method in suppressing the scale dependence compared to the results for the OPT. An important aspect elucidated by these results is the difference in scale at the minimum of the potential, a feature that had already been demonstrated in Fig. 6(a).
VI Conclusions
@@
In this work, we present an alternative resummation method that aims to improve the effective potential obtained with the traditional OPT by imposing RG conditions. The new prescription, introduced as the Variational Renormalization Group (VRG), combines the Renormalization Group Improvement (RGI) method, originally prescribed in Refs. [68, 69], applied to the effective potential of the Optimized Perturbation Theory (OPT) for a real scalar field model at finite temperatures. This approach encapsulates a consistent way to merge these two tools and the results presented here for the $\lambda\phi^{4}$ theory are promising, showing a significant reduction in the scale dependence in the symmetric phase, especially for coupling values such that $\lambda\gtrsim 1$. A major advantage of the prescription proposed in this work is that no modification of the standard OPT framework is required apart from the incorporation of some properties of the RGI procedure. Since VRG improves the scale dependence of the standard OPT we believe that it can also be useful in improving the scale other thermal resummation methods, such as SPT and HTLpt, which are plagued by high scale dependence issues.
@@
In the broken phase, the VRG has proven to be an excellent tool for predicting investigating phase transition patterns. In particular, the predictions for the order parameter and the critical temperature have proven to be very stable against scale variations. In this application, it was not possible to access the critical temperature for higher couplings because of the use of high-temperature approximations, which may impose limitations on the results for high couplings. In this case, the safest approach would be to perform a more comprehensive numerical analysis to better understand the broken phase for a wide range of $\lambda$ values. Finally, we have observed that the VRG respects the same universality class as the $\lambda\phi^{4}$ theory, predicting a second-order phase transition at $\delta^{2} order$, as expected.
@@ Acknowledgements.
The authors acknowledge financial support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) - Finance Code 001. R.O.R. is also partially supported by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. M.B.P. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Process No. 307261/2021-2 and 403016/2024-0.
@@ Appendix B Renormalized Parameters
In the context of the usual perturbation theory, the renormalized parameters are determined by the renormalization conditions (e.g. by the pole of the propagator for the mass and by the amplitude $2\to 2$ scattering for the quartic coupling constant) [81]. These parameters are also known as physical parameters [80]. Accordingly, the bare mass and coupling are related to the reormalized ones through
$m_{b}^{2}$ $=$ $m^{2}+\frac{\hbar\lambda m^{2}}{2(4\pi)^{2}}\left[\ln\left(\frac{\mu^{2}}{m^{2}}\right)+1\right]+\frac{\hbar^{2}\lambda^{2}m^{2}}{(4\pi)^{4}}\left[-\frac{11}{48}\right$. (48)
@@
$\lambda_{b}$ $=$ $\lambda+\frac{\hbar\lambda^{2}}{(4\pi)^{2}}\left[\frac{3}{2}\ln\left(\frac{\mu^{2}}{m^{2}}\right)+1\right]+\mathcal{O}(\lambda^{3}).$ (49)
@@
${\cal F}_{\delta}$ $=$ $\delta\frac{\hbar\lambda\varphi^{2}\Omega^{2}}{4(4\pi)^{2}}\left[L_{\Omega}+1\right]+\delta\frac{\hbar^{2}\lambda\Omega^{2}}{4(4\pi)^{4}}\left[L_{\Omega}+1\right]$ (50)
@@
${\cal F}_{\delta^{2}}$ $=$ $\delta^{2}\left\{\frac{\hbar^{2}\eta^{2}\lambda}{4(4\pi)^{4}}\left[J_{2,\Omega}(L_{\Omega}+1)\Omega^{2}-J_{1,\Omega}T^{2}(2J_{2,\Omega}+L_{\Omega})\right]\right$. (51)
@@
Appendix C Solutions for the RGI functions
@@
The solutions for the scale dependent mass at orders $\hbar^{0}$, $\hbar$ and $\hbar^{2}$ are found similarly (using the boundary conditions $\bar{m}_{0}^{2}(0)=m^{2}(\mu)$ and $\bar{m}_{1}^{2}(0)=\bar{m}_{2}^{2}(0)=0$) and are, respectivley, given by
@@
Finally, the solutions for the vacuum energy at orders $\hbar^{0}$, $\hbar$ and $\hbar^{2}$, using the boundary conditions $\bar{\Lambda}_{0}(0)=\Lambda(\mu)$ and $\bar{\Lambda}_{1}(0)=\bar{\Lambda}_{2}(0)=0$, are found to be given, respectively, by
$\bar{\Lambda}_{0}$ $=$ $\Lambda-\frac{m^{4}\beta_{\Lambda 0}}{\lambda(\beta_{0}-2\gamma_{m0})}[\xi^{1-2\gamma_{m0}/\beta_{0}}-1],$ (67)
@@
Appendix D VRG Effective Potential
@@
$V_{\rm VRG}^{\delta}$ $=$ $\Lambda+\frac{1}{2}\bar{\Omega}^{2}\varphi^{2}+\frac{1}{4!}\frac{\lambda}{\xi}\varphi^{4}-\frac{\bar{\Omega}^{4}\left(2L_{\bar{\mu}}+3\right)}{8\left(4\pi\right)^{2}}$ (70)
@@
where $\xi=1-\lambda\beta_{0}t$, $t=\ln(\bar{\mu}/\mu)$, $\bar{\Omega}^{2}=(\frac{m^{2}}{\xi^{1/3}}+\eta^{2})$, $L_{\bar{\mu}}=\ln(\mu^{2}e^{2t}/\bar{\Omega}^{2})$, and
$F_{\rm vac,1}$ $=$ $-\frac{m^{4}}{2\lambda}\left(\xi^{1/3}-1\right)+\frac{m^{4}}{(4\pi)^{2}}\left[\frac{35}{54}\left(\frac{1}{\xi^{2/3}}-1\right)+\frac{17\ln\xi}{54\xi^{2/3}}+\frac{19}{54}\left(\xi^{1/3}-1\right)\right]$ (71)
$+$ $\frac{m^{2}\lambda}{108\xi^{4/3}(4\pi)^{4}}(19(\xi-1)-34\ln\xi)\left\{\left[\left(L_{\bar{\mu}}+1\right)\bar{\Omega}^{2}-J_{1,\bar{\Omega}}T^{2}\right]-\eta^{2}\left[L_{\bar{\mu}}+J_{2,\Omega}\right]\right\}$
- $\frac{m^{2}\lambda\varphi^{2}(8(\xi-1)-17\ln\xi)}{54\xi^{4/3}(4\pi)^{2}}-\frac{m^{4}\lambda}{\xi^{5/3}(4\pi)^{4}}\left[-\frac{289\ln^{2}\xi}{1458}+\frac{323(\xi-1)\ln\xi}{1458}\right$.
- $\frac{36\zeta(3)+23}{30}\xi^{5/3}+\frac{7776\zeta(3)+2509}{11664}\xi^{2}-\left.\frac{3888\zeta(3)+9239}{29160}+\frac{7776\zeta(3)+10129}{11664}\xi\right].$
@@
$+$ $7L_{\bar{\mu}}^{2}+\left(\frac{57}{2}+4C_{1}+\pi^{2}\right)L_{\bar{\mu}}-5-\frac{11\pi^{2}}{12}-\left.\psi^{\prime\prime}\left(1\right)+C_{0}+4C_{1}\right)\bar{\Omega}^{4}+\left(2J_{1,\bar{\Omega}}^{2}+J_{1,\bar{\Omega}}^{2}J_{2,\bar{\Omega}}\right$.

### v2

@@ IV Results
Let us now analyze how well the VRG performs in order to attenuate the scale dependence of the OPT effective potential by mainly comparing predictions for the symmetric ($m^{2}\geq 0$) and the non-symmetric ($m^{2}<0$) cases. Concerning the first case, we will also make a comparison with the results provided by other nonperturbative methods commonly used in the literature, e.g., the results obtained from the 2-particle irreducible approach (2PI), from the functional renormalization group (FRG) and from the RGOPT method. This will help us gauge how our approach performs compared to these other methods. In addition, we also examine whether the VRG potential preserves the universality class of the $\lambda\phi^{4}$ theory [^1] Recall that the $Z_{2}$ symmetric scalar field model (1) belongs to the same universality class of the Ising model for $d\leq 4$, with the model exhibiting a second-order phase transition at the critical point [70, 59, 71]..
@@
To analyze the dependence of the effective potential on the renormalization scale, we start by considering the scale dependence of the renormalized parameters $m(\mu)^{2}$, $\lambda(\mu)$, and $\varphi(\mu)$. As already specified in the previous section, at the first order for both OPT and VRG, we consider the running of the renormalized parameters up to ${\cal O}(\hbar^{2})$ and which are determined by the equations:
@@
where we now substitute these solutions in Eq. (6), up to order $\delta^{2}$, in the case of OPT, and in Eq. (78) for VRG. The boundary conditions used to solve the set of Eqs. (27)-(32) are: $\lambda(\mu_{0})=\lambda_{0}$, $m(\mu_{0})=m_{0}$, and $\varphi(\mu_{0})=\varphi_{0}$ where $\mu_{0}$ is a reference scale to be defined below. As is common in the literature, to check the stability of the computed quantities with the scale $\mu$, we will vary it around the so-called “central” value $\mu=2\pi T$ within the usual range $\mu\in[\pi T,4\pi T]$. As a reference scale, we can then choose $\mu_{0}=2\pi T_{0}$ where, for simplicity, the reference temperature is set to $T_{0}=m_{0}/(2\pi)$ so that $\mu_{0}=m_{0}$. This choice will allow us to easily express all physical quantities in units of $m_{0}$ (or likewise, in terms of the reference scale $\mu_{0}$).
IV.1 Symmetric Phase
@@ IV.2 Broken Phase
As usual, the order parameter $\sigma(T)$ can be determined by minimizing the thermal effective potential, which here is done for both the OPT and VRG ones,
@@
Figure 8: Subtracted effective potential, $\Delta V=V_{\rm eff}(\varphi,T)-V_{\rm eff}(\varphi=0,T)$, as a function of $\varphi$. The results for OPT and VRG are at order $\delta^{2}$. The coupling is fixed at the values $\lambda_{0}=1$. The scale dependence range is such that $\pi T\leq\mu\leq 4\pi T$. From bottom to top, the regions correspond to the temperature fixed at $T_{c}-\Delta T, T=T$$T_{c}$, and $T=T_{c}+\Delta T$, respectively, with $\Delta T/m_{0}=5\times 10^{-6}$.
@@
Let us now check the dependence of the effective potential in the OPT and VRG cases as a function of the background field just as we have done in the symmetric case (see Fig. 5). This is shown in Fig. 8. We choose $\lambda_{0}=1$ as a representative value for the coupling constant. The values of the temperature have been taken below and above the critical value, which shows well that the phase transition is second-order, as also confirmed with the results shown in Fig. 6. Furthermore, the results displayed in Fig. 8 indicate the efficiency of the VRG method in suppressing the scale dependence compared to the results for the OPT. An important aspect elucidated by these results is the difference in scale at the minimum of the potential, a feature that had already been demonstrated in Fig. 6(a).
V Conclusions
In this work, we present an alternative resummation method that aims to improve the effective potential obtained with the traditional OPT by imposing RG conditions. The proposed new prescription, called here the Variational Renormalization Group (VRG), combines the Renormalization Group Improvement (RGI) method, originally prescribed in Refs. [69, 58] with the OPT procedure. The OPT is directly applied to the RGI finite-temperature effective potential, which then combines the variational resummation with a systematic reduction of renormalization-scale dependence.
@@ V Conclusions
This approach encapsulates a consistent way to merge these two tools and the results presented here for the $\lambda\phi^{4}$ theory are promising, showing a significant reduction in the scale dependence in both the symmetric and broken phases of the theory. A major advantage of the prescription proposed in this work is that no modification of the standard OPT framework is required apart from the incorporation of the properties of the RGI procedure. Since VRG improves the scale dependence of the standard OPT, we believe that it can also be useful in improving the scale of other thermal resummation methods, such as SPT and HTLpt, which are plagued by high scale dependence issues.
@@
In the broken phase, the VRG has proven to be fully consistent for what we expect for the phase transition pattern of the theory. In particular, the predictions for the order parameter and the critical temperature have proven to be very stable against scale variations. We have observed that the VRG respects the same universality class as the $\lambda\phi^{4}$ theory, predicting a second-order phase transition, as expected. In this application, it was not possible to access the critical temperature for higher couplings because of the use of high-temperature approximations. This may be remediated in the future by accounting for the full momentum integration of the loop terms, bypassing the high-temperature approximations. This would require to perform a more comprehensive numerical analysis to better understand the broken phase for a wide range of $\lambda$ and temperature values. We leave this more involved numerical analysis for a future work.
@@ Acknowledgements.
The authors thank Jean-Loïc Kneur for discussions concerning the method considered here. The authors acknowledge financial support of the Coordenação de Aperfeiçoamento de Pessoal de Nível Superior (CAPES) - Finance Code 001. R.O.R. is also partially supported by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021. M.B.P. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Process No. 307261/2021-2 and 403016/2024-0.
@@ Appendix B Renormalized Parameters
In the context of the usual perturbation theory, the renormalized parameters are determined by the renormalization conditions (e.g. by the pole of the propagator for the mass and by the amplitude $2\to 2$ scattering for the quartic coupling constant) [81]. These parameters are also known as physical parameters [80]. Accordingly, the bare mass and coupling are related to the renormalized ones through
$m_{b}^{2}$ $=$ $m^{2}+\frac{\lambda m^{2}}{2(4\pi)^{2}}\left[\ln\left(\frac{\mu^{2}}{m^{2}}\right)+1\right]+\frac{\lambda^{2}m^{2}}{(4\pi)^{4}}\left[-\frac{11}{48}\right$. (47)
@@
$\lambda_{b}$ $=$ $\lambda+\frac{\lambda^{2}}{(4\pi)^{2}}\left[\frac{3}{2}\ln\left(\frac{\mu^{2}}{m^{2}}\right)+1\right]+\mathcal{O}(\lambda^{3}).$ (48)
@@
${\cal F}_{\delta}$ $=$ $\delta\frac{\lambda\varphi^{2}\Omega^{2}}{4(4\pi)^{2}}\left[L_{\Omega}+1\right]+\delta\frac{\lambda\Omega^{2}}{4(4\pi)^{4}}\left[L_{\Omega}+1\right]$ (49)
@@
[…]
- $\left.\left.J_{1,\Omega}T^{2}(2J_{2,\Omega}+L_{\Omega})\right]\right$.
@@
Appendix C The perturbative effective potential of the massive $\lambda\phi^{4}$ theory in $\hbar$ order
In this appendix, we present the effective potential of the massive $\lambda\phi^{4}$ theory at finite temperature, organized in orders of $\hbar$. The zero temperature case was used at the beginning of the section III for the review of RGI, whose result is found in Ref. [58]
$V_{\rm eff}$ $=$ $\Delta V_{\rm eff}^{\hbar^{0}}\hbar^{0}+\Delta V_{\rm eff}^{\hbar^{1}}\hbar^{1}+\Delta V_{\rm eff}^{\hbar^{2}}\hbar^{2}$ (51)
$+$ $\Delta V_{\rm eff}^{\hbar^{3}}\hbar^{3}+{\cal O}(\hbar^{4}),$
where
$\Delta V_{\rm eff}^{\hbar^{0}}$ $=$ $\frac{1}{2}m^{2}\varphi^{2}+\frac{\lambda}{4!}\varphi^{4}+\Lambda,$ (52)
$\Delta V_{\rm eff}^{\hbar^{1}}$ $=$ $-\frac{m_{H}^{4}\left(2L_{m_{H}}+3\right)}{8\left(4\pi\right)^{2}}-\frac{J_{0,m_{H}}T^{4}}{2\left(4\pi\right)^{2}},$ (53)
$\Delta V_{\rm eff}^{\hbar^{2}}$ $=$ $\frac{\lambda}{8\left(4\pi\right)^{4}}\left[\left(L_{m_{H}}+1\right)m_{H}^{2}-J_{1,m_{H}}T^{2}\right]^{2}$ (54)
- $\frac{\lambda^{2}\varphi^{2}}{12(4\pi)^{4}}\left[-3m_{H}^{2}\left(L_{m_{H}}^{2}+3L_{m_{H}}+C_{2}\right)\right$.
$+$ $\left.3T^{2}\left(L_{m_{H}}J_{1,m_{H}}+H_{2,m_{H}}+H_{3,m_{H}}\right)\right]$
- $\frac{\lambda^{2}\varphi^{2}m_{H}^{2}}{8(4\pi)^{4}}\left[(L_{m_{H}}+1)^{2}+1+\frac{\pi^{2}}{6}\right]\;,$
and
$\Delta V_{\rm eff}^{\hbar^{3}}$ $=$ $\frac{\lambda^{2}}{48\left(4\pi\right)^{6}}\left[\left(\left(3L_{m_{H}}+4\right)J_{1,m_{H}}^{2}+J_{1,m_{H}}^{2}J_{2,m_{H}}+2K_{2,m_{H}}+\frac{4}{3}K_{3,m_{H}}\right)3T^{4}\right$. (55)
- $\left(12L_{m_{H}}^{2}+28L_{m_{H}}-12-\pi^{2}-4C_{0}+6\left(L_{m_{H}}+1\right)J_{2,m_{H}}\right)J_{1,m_{H}}m_{H}^{2}T^{2}$
$+$ $\left.\left(5L_{m_{H}}^{3}+17L_{m_{H}}^{2}+\frac{41}{2}L_{m_{H}}-23-\frac{23\pi^{2}}{12}-\psi^{\prime\prime}\left(1\right)+C_{1}+3\left(L_{m_{H}}+1\right)^{2}J_{2,m_{H}}\right)m_{H}^{4}\right]+{\cal O}(\lambda^{3})\;,$
where we have defined $L_{m_{H}}=\ln(\mu^{2}/m_{H}^{2})$ and $m_{H}^{2}=m^{2}+\frac{1}{2}\lambda\varphi^{2}$. Terms beyond ${\cal O}(\lambda^{2})$ will not be necessary for the purposes of this work. This potential contains all orders in perturbation theory in $\lambda$. The quantities $J_{0,m_{H}}$, $J_{1,m_{H}}$, $J_{2,m_{H}}$, $H_{2,m_{H}}$, $H_{3,m_{H}}$, $K_{2,m_{H}}$ and $K_{3,m_{H}}$ are all functions of $m_{H}/T$ and their explicit expressions are given in App. A.
Appendix D Solutions for the RGI functions
@@
The solutions for the scale dependent mass at orders $\hbar^{0}$, $\hbar$ and $\hbar^{2}$ are found similarly (using the boundary conditions $\bar{m}_{0}^{2}(0)=m^{2}(\mu)$ and $\bar{m}_{1}^{2}(0)=\bar{m}_{2}^{2}(0)=0$) and are, respectively, given by
@@
Finally, the solutions for the vacuum energy at orders $\hbar^{0}$, $\hbar$ and $\hbar^{2}$, using the boundary conditions $\bar{\Lambda}_{0}(0)=\bar{\Lambda}_{1}(0)=\bar{\Lambda}_{2}(0)=0$, are found to be given, respectively, by
$\bar{\Lambda}_{0}$ $=$ $-\frac{m^{4}\beta_{\Lambda 0}}{\lambda(\beta_{0}-2\gamma_{m0})}[\xi^{1-2\gamma_{m0}/\beta_{0}}-1],$ (71)
@@
Appendix E VRG Effective Potential
@@
$V_{\rm VRG}^{\delta}$ $=$ $\frac{1}{2}\bar{\Omega}^{2}\varphi^{2}+\frac{1}{4!}\frac{\lambda}{\xi}\varphi^{4}-\frac{\bar{\Omega}^{4}\left(2L_{\bar{\mu}}+3\right)}{8\left(4\pi\right)^{2}}$ (74)
@@
where $\xi=1-\lambda\beta_{0}t$, $t=\ln(\bar{\mu}/\mu)$ and we have also defined
$\bar{\Omega}^{2}$ $=$ $\frac{1}{\xi^{1/3}}(m^{2}+\eta^{2}),$ (75)
$L_{\bar{\mu}}$ $=$ $\ln(\mu^{2}e^{2t}/\bar{\Omega}^{2}),$ (76)
and
$F_{\rm vac,1}$ $=$ $\frac{\left(\eta^{2}-\xi^{1/3}\bar{\Omega}^{2}\right)^{2}}{2\lambda}\left(1-\xi^{1/3}\right)+\frac{\bar{\Omega}^{4}}{54(4\pi)^{2}}\left[35-54\xi^{2/3}+17\ln\xi+19\xi\right]$
$+$ $\frac{\bar{\Omega}^{2}\lambda}{108\xi(4\pi)^{4}}\left[19(\xi-1)-34\ln\xi\right]\left[\left(L_{\bar{\mu}}+1\right)\bar{\Omega}^{2}-J_{1,\bar{\Omega}}T^{2}\right]$
- $\frac{\bar{\Omega}^{2}\lambda\varphi^{2}\left[8(\xi-1)-17\ln\xi\right]}{54\xi(4\pi)^{2}}-\frac{\bar{\Omega}^{4}\lambda}{\xi(4\pi)^{4}}\left[-\frac{289\ln^{2}\xi}{1458}+\frac{7776\zeta(3)+2509}{11664}\xi^{2}\right$.
$+$ $\left.\frac{323(\xi-1)\ln\xi}{1458}-\frac{36\zeta(3)+23}{30}\xi^{5/3}-\frac{3888\zeta(3)+9239}{29160}+\frac{7776\zeta(3)+10129}{11664}\xi\right].$
@@
$V_{\rm VRG}^{\delta^{2}}$ $=$ $V_{\rm VRG}^{\delta}+\frac{\eta^{4}}{4\xi^{2/3}(4\pi)^{2}}\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)+\frac{\lambda\eta^{2}}{4\xi^{4/3}\left(4\pi\right)^{4}}\left[J_{1,\bar{\Omega}}J_{2,\bar{\Omega}}T^{2}+L_{\bar{\mu}}\left(L_{\bar{\mu}}+1\right)\bar{\Omega}^{2}\right]$ (78)
- $\frac{\lambda^{2}}{48\xi^{2}\left(4\pi\right)^{6}}\left[\left(5L_{\bar{\mu}}^{3}+7L_{\bar{\mu}}^{2}+\left(\frac{57}{2}+4C_{1}+\pi^{2}\right)L_{\bar{\mu}}-5-\frac{11\pi^{2}}{12}-\psi^{\prime\prime}\left(1\right)+C_{0}+4C_{1}\right)\bar{\Omega}^{4}\right$.
$+$ $\left.\left(2J_{1,\bar{\Omega}}^{2}+J_{1,\bar{\Omega}}^{2}J_{2,\bar{\Omega}}+2K_{2,\bar{\Omega}}+\frac{4}{3}K_{3,\bar{\Omega}}\right)3T^{4}\right]$
@@
$+$ $\left.\left.\left(\frac{C_{1}}{3}+C_{2}+\frac{L_{\bar{\mu}}}{6}\right)\bar{\Omega}^{2}\right]\right\}+\frac{\lambda^{2}\varphi^{4}}{48\xi^{2}(4\pi)^{2}}(2-3J_{2,\bar{\Omega}})+F_{\rm vac,2}+F_{\varphi},$
where
$F_{\varphi}$ $=$ $\varphi^{2}\left\{\frac{\lambda^{2}\bar{\Omega}^{2}(19(\xi-1)-34\ln\xi)}{216\xi^{2}(4\pi)^{4}}\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)-\frac{\lambda^{2}(\xi+34\ln\xi-1)}{72\xi^{2}(4\pi)^{4}}\left[(L_{\bar{\mu}}+1)\bar{\Omega}^{2}-T^{2}J_{1,\bar{\Omega}}\right]\right$.
$+$ $\left.\frac{\lambda^{2}\bar{\Omega}^{2}}{\xi^{2}(4\pi)^{4}}\left[\frac{17(43+108L_{\bar{\mu}}-16\xi)\ln\xi}{5832}+\frac{289\ln^{2}\xi}{729}+\frac{(\xi-1)}{11664}(-864L_{\bar{\mu}}+973\xi+3888\zeta(3)(\xi-1)-4186)\right]\right$.
$+$ $\left.\frac{\eta^{2}\lambda(8(\xi-1)-17\ln\xi)}{54\xi^{4/3}(4\pi)^{2}}\right\}+\frac{\lambda^{2}\varphi^{4}\left(\xi-1+17\ln\xi\right)}{216\xi^{2}(4\pi)^{2}},$
and
$F_{\rm vac,2}$ $=$ $\frac{\eta^{4}\left(35-54\xi^{2/3}+19\xi+17\ln\xi\right)}{54(4\pi)^{2}\,\xi^{2/3}}+\frac{17\lambda^{2}\ln\xi}{72\xi^{2}(4\pi)^{6}}\left[(L_{\bar{\mu}}+1)\bar{\Omega}^{2}-T^{2}J_{1,\bar{\Omega}}\right]^{2}+\frac{\eta^{2}\lambda\bar{\Omega}^{2}}{\xi^{4/3}(4\pi)^{4}}\left[-\frac{23\xi^{5/3}}{15}+\frac{2509\xi^{2}}{5832}\right$. (80)
$+$ $\left.\frac{10129\xi}{5832}-\frac{289\ln^{2}\xi}{729}+\frac{323}{729}\xi\ln\xi-\frac{323\ln\xi}{729}-\frac{1}{5}12\xi^{5/3}\zeta(3)+\frac{4\xi^{2}\zeta(3)}{3}+\frac{4\xi\zeta(3)}{3}-\frac{4\zeta(3)}{15}-\frac{9239}{14580}\right]$
$+$ $\frac{\lambda^{2}\bar{\Omega}^{2}(19(\xi-1)-34\ln\xi)^{2}}{23328\xi^{2}(4\pi)^{6}}\left\{\left[(L_{\bar{\mu}}+1)\Omega^{2}-T^{2}J_{1,\bar{\Omega}}\right]-2\bar{\Omega}^{2}\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)\right\}+\left\{\left[(L_{\bar{\mu}}+1)\Omega^{2}-T^{2}J_{1,\bar{\Omega}}\right]\right$.
- $\left.2\bar{\Omega}^{2}\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)\right\}\frac{\lambda^{2}(L_{\bar{\mu}}+1)\bar{\Omega}^{2}(19(\xi-1)-34\ln\xi)}{432\xi^{2}(4\pi)^{6}}-\frac{\lambda^{2}\bar{\Omega}^{2}(19(\xi-1)-34\ln(\xi))}{216\xi^{2}(4\pi)^{6}}\left(L_{\bar{\mu}}+J_{2,\bar{\Omega}}\right)\times$